A body of mass 2kg moves 30m aginst a frictional force of 100N. If the work is completely converted into heat, the amount of heat expressed in joules will be?

Grade:10

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Vishnu

The work done against friction will be  converted into heat

Work done against friction=100 x 30 = 3000 J

Heat=3000 J Answer

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

force =100N

workdone by friction = f.s = force*displacement

                             =100*30 =3*103 J

this is the energy which is converted to heat ...